A boat moving downstream (with the current) in a river takes 4 min and 7secs to cover

a measured mile according to marks on shore, and 7 min and 30 secs to cover the same

distance moving upstream (against the current). If the boat’s velocity relative to the

water is 11 km/h, what is the velocity of the current in the river?


1
Expert's answer
2021-10-14T11:33:52-0400

Express the time of motion downstream and upstream:


"t_d=\frac{d}{v+u},\\\space\\
t_u=\frac{d}{v-u},\\\space\\
t_dv+t_du=t_uv-t_uu,\\\space\\
u=v\frac{t_d-t_u}{t_d+t_u}=1.24\text{ km/h}."